Y
re 6, Video e Z e7 Video 8 Video Materials al videos ts and due Remaining Time: 58 minutes, 21 seconds. Question Completion Status: Question 1 of 10 1 points Save Answer You have just won a lottery, which offers a choice in how you may receive compounded annually, whic your prize. If the interest rate is 10% per annum h of the following prizes has the highest present value? $2 000 000 immediately O An ordinary annuity of $300 000 per year for each of the next 10 years (starting one year from now) O $3 000 000 at the end of five years perpetuity of $250 000 per year starting one year from now to the next question prevents changes to this answer Question 1 of 10
Answer #1
A B C D E F G
2
3 Interest rate 10%
4
5 $2000,000 paid immediately will have present value of $2,000,000.
6
7 Present value annuity of $300,000 per year for next 10 year. =300,000*(P/A,10%,10)
8 $1,843,370.13 =300000*PV(D3,10,-1,0)
9
10 Present value of $3,000,000 received at the end of 5 years =3,000,000*(P/F,10%,5)
11 $1,862,763.97 =3000000*(1/((1+D3)^5))
12
13 Present value of perpetuity of $250,000 per year =$250,000 / 10%
14 $2,500,000.00 =250000/D3
15
16 Thus the present value of perpetuity of $250,000 is the highest,
17 therefore the fourth option is correct.
18

Formula sheet

A B C D E F G
2
3 Interest rate 0.1
4
5 $2000,000 paid immediately will have present value of $2,000,000.
6
7 Present value annuity of $300,000 per year for next 10 year. =300,000*(P/A,10%,10)
8 =300000*PV(D3,10,-1,0) =300000*PV(D3,10,-1,0)
9
10 Present value of $3,000,000 received at the end of 5 years =3,000,000*(P/F,10%,5)
11 =3000000*(1/((1+D3)^5)) =3000000*(1/((1+D3)^5))
12
13 Present value of perpetuity of $250,000 per year =$250,000 / 10%
14 =250000/D3 =250000/D3
15
16 Thus the present value of a perpetuity of $250,000 is the highest,
17 therefore the fourth option is correct.
18